A 0.56-C charge is placed 15.0 m from a 1.56-C charge. What is the force between them?
1
Expert's answer
2019-04-15T10:06:27-0400

The force between them:


"F=\frac{kq_1q_2}{r^2}"

"F=\frac{(8.99 \cdot 10^9)(0.56)(1.56)}{15^2}=3.49 \cdot 10^7 N"